Kimberly is twice as old as ryan. in 4 years she will be 7 years older than ryan. how old is kimberly now?
Answer by ankor@dixie-net.com(22740)

Kimberly is twice as old as ryan.
k = 2r
:
in 4 years she will be 7 years older than ryan.
:
This statement is made to confuse, K is 7 yrs older than R now, in 4 years, and until death,
Just say
k = r + 7
replace k with 2r from the 1st statement
2r = r + 7
2r - r = 7
r = 7 is R's present age
and obviously
2(7) = 14 is Kimberly's present age